OpenStudy (hpfan101):

Ultraviolet radiation falls in the wavelength region of 1.00×10-8 to 1.00×10-7 meters. What is the wavelength of ultraviolet radiation that has an energy of 1.99×10-20 kJ?

OpenStudy (michele_laino):

the requested wavelength \(\lambda\), is given by the subsequent formula: \[\Large \lambda = \frac{{hc}}{\mathcal{E}}\] where \(h\) is the constant of Planck, \(c\) is the speed of light, and: \(\Large \mathcal{E}=1.99 \cdot 10^{-17}\) joules

OpenStudy (hpfan101):

So after doing this, would our answer be \[9.982 \times 10^{-9}\] ?

OpenStudy (hpfan101):

Units would be meters